OpenStudy (anonymous):

PLEASE HELP! Algebra 1 Standard, 9th Grade y-4 divided by 2 =10

OpenStudy (anonymous):

(y - 4 ) / 2 = 10. First, expand into two fractions as y/2 - 4/2 = 10 then simply as y/2 - 2 = 10 then add 2 to both sides to get y/2 = 8 then multiply bot sides by 2 to get y = 16

OpenStudy (anonymous):

THANK YOU!!!!!